Journalist Report – January 13th



Crew 272 Journalist Report 1-13-2023

Sol: 12

Author: Kenny Pritchard, Crew Journalist

Title: Lakshya (“Aim”)

For our final crew breakfast on Mars, despite rumors of lasagna, Adriana whipped up some biscuits and rehydrated butter. Our EVA this morning was planned for an early 9:00 a.m. departure, so we had little time to dilly dally – we ran an oxymoronically compacted stretching session, and four of us suited up for the EVA with biscuits still in our mouths.

The mission today was a real search and rescue to find and retrieve the Boiler Transmission Station (BTS) which had been deployed to a ledge near the Hab (we deemed it Boiler Ledge) on a previous EVA. After a few powerful dust and wind storms, the crew was concerned that it may have been blown far away from its last known location and lost to the sands of Mars. Mason, Arly, Megan, and I took the rovers to a good stopping point, walked up a familiar hill, and began scouting toward the expected location with the drone. This cycle of walking and scouting repeated a few times until we got close enough to where the drone camera could almost make out our target and we could walk over to Boiler Ledge. Lo and behold, the BTS sat serenely waiting our arrival, if a little waterlogged, and missing its original housing box. We collected the treasured transmitter and took some good drone videos and pictures hoisting our Purdue “Boiler Up” flag near the edge of Boiler Ledge. Having made excellent time to the Ledge and finding the objective easily, we made it back to the Hab before noon, where the rest of the crew sat working on the final tasks of our rotation: finishing up research, cleaning up, packing, and a writing a whole slew of end-of-mission reports.

With our time at MDRS coming to an end, it’s a good time to reflect on the progress we made. Although our crew had been meeting and planning for months leading up to our rotation, we had only just seen the tip of the iceberg of each other’s personalities and life stories (and flaws) (just kidding). Living together in an enclosed space on an inhospitable planet for two weeks was a quick fix! We’ve voiced our pet peeves, become adept with our equipment, endured exhausting EVAs, overcome unexpected challenges, been sim-resurrected from several sim-deaths, formed alliances and fought wars over board games, and we can proudly say that every crew member has achieved at least the minimum objectives of their research. The difference between our first steps and what we’re capable of accomplishing now is night and sol.

One common theme of each sol has been abundant references to the 2004 Bollywood movie Lakshya, shared with us by our fearless commander Kshitij. In this film, the main character asks himself “Main Aisa Kyun Hoon” – “Why am I like this?” and struggles to find his aim in life. I’d compare our first meetings and our final sols here with the beginning and end of the film, our crew taking a dynamic arc from inexperience and lack of direction to a resolute sense of purpose and efficiency, thanks largely to Kshitij’s wisdom and leadership (and humor). It’s been a good run here on Mars, but today we end our simulation to prepare for the long ride back to Earth! On behalf of the entire crew, thank you sincerely to all the loved ones who followed our journey and kept in touch along the way. Your support means… the world!

Signing off from Mars,

Adriana “Rocky” Brown, Arly “Maple” Black, Kenny “Hemingway” Pritchard, Dr. Kshitij “Chai” Mall, Madelyn “Sprout” Whitaker, Mason “Scout” Kuhn, and Megan “Xerox” Rush

Operations Report – January 13th



Crew 272 Operations Report 13-1-2023

SOL: 12

Name of person filing report: Mason Kuhn

Non-nominal systems: Oven light

Notes on non-nominal systems: No additional issues


Spirit rover used: Yes

Hours: 207.3

Beginning charge: 100%

Ending charge: 73%

Currently charging: Yes

Opportunity rover used: Yes

Hours: 112.0

Beginning charge: 100%

Ending charge: 78%

Currently charging: Yes

Curiosity rover used: No

Hours: 219.3

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Perseverance rover used: No

Hours: 253.9

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

General notes on rovers: Spirit and Opportunity were used on the EVA. They performed nominally. The chargers were draped over the front bumper upon return to the Hab. Perseverance is still at the outpost temporarily.

Summary of Hab operations: Post-sim cleanings to be conducted in the afternoon (upper deck, RAM, Sci Dome, GreenHab), and cleanings of the lower deck will be completed this afternoon with touch-up cleanings taking place in the morning of Sol 13.

WATER USE: 11.8 gallons (by 2:30pm and before cleanings that use water)

Water (static tank): 142.8 gallons

Static tank pipe heater (on or off): On

Static tank heater (On or off): On

Toilet tank emptied: No

Summary of internet: Nominal

Summary of suits and radios: Suits 1, 5, 9, and 10 were used on the EVA. All suits performed nominally, and all are cleaned & currently charging. All radios have been charged to full, and they have been removed from the charging docks. The EVA crew had no communication issues.

Summary of GreenHab operations: Monitored and watered GreenHab crops, with a bit of extra watering due to higher temperatures. Fertilized in the afternoon. Harvested beans, chives, thyme, and dill. Planted cilantro and radishes. Cleared out dead leaves and overgrowth, distributed organic matter around GreenHab to recycle nutrients, and swept and cleaned the floor & cabinets.

WATER USE: 9 gallons

Heater: On

Supplemental light: 4 hours

Harvest: 15g beans, 3g chives, 15g thyme, 30g dill

Summary of Science Dome operations: Crew items were removed from the Science Dome, and cleanings according to the checklist to be completed shortly

Dual split: On (from approximately 8:00 pm to 8:00 am)

Summary of RAM operations: All tools from the RAM that were in the lower deck have been returned to the RAM after final measurements were made, cleanings were conducted.

Summary of any observatory issues: N/A

Summary of health and safety issues: N/A

Questions, concerns, and requests to Mission Support: N/A

GreenHab Report – January 13th



Crew 272 GreenHab Report 13-01-2023

GreenHab Officer: Madelyn Whitaker

Environmental control: heater

Average temperatures: 78 F

Hours of supplemental light: 4 hours

Daily water usage for crops: 9

Daily water usage for research and/or other purposes: 0

Water in Blue Tank : 212 gallons

Time(s) of watering for crops: 1000, 1400

Changes to crops: Planted cilantro and radish. Harvested chives, dill, thyme, beans.

Narrative: Monitored and watered GreenHab crops, temperatures were very warm so I watered a bit extra. Fertilized in the afternoon. Harvested beans, chives, thyme, dill. Planted cilantro and radishes. Cleared out some dead leaves and overgrowth, distributed organic matter around GreenHab to recycle nutrients. Swept and cleaned floor and cabinets.

Harvest: (include which crop and mass in grams): Harvested 15g beans, 3g chives, 15g thyme, 30g dill

Support/supplies needed: N/A

EVA Report – January 13th



Crew 272 EVA Report 13-01-2023

EVA # 10

Author: Arly Black

Purpose of EVA: Simple simulation of search and rescue of an incapacitated astronaut. In this case, the astronaut was the Boiler Transmission Station (BTS) located at N519600, E4251500.

Start time: 9:00 AM

End time: 10:50 AM

Narrative: Today’s EVA started off a lot earlier than the rest of them have this mission. This was the last EVA of our rotation (noooo!) and we were eager to get started and save our lost astronaut/Kshitij’s baby (and also to get back early to write our many final mission reports). After wiping the sleep from our eyes, a quick stretch, and a rushed breakfast, the EVA crew jumped into our suits, which by now we need no assistance to put on. Five minutes of space jokes (Why aren’t astronauts hungry when they get to space? Because they had a big launch!) made the airlock prebreathing pass by quickly. It was a chilly morning and when we got to our rovers, there was frost on our seats. Some of us wish we had noticed that before sitting down… Arly and Mason led the way to our destination in Spirit (and with spirit!) with Kenny and Megan following in Opportunity. We drove to the Gateway to Candor and parked, using the mound on the right side of the road to help us navigate on-foot to our destination, based on experience from the last EVA to the self-named Boiler Ledge. When we reached the top of the hill, Mason flew our search and rescue drone to see if our lost astronaut BTS (transmitter + box, not the K-Pop band) was still there. He was unable to see anything and lost connection at some point, so we continued walking towards the ledge. At a closer point about 300 ft from the ledge, Mason again attempted to spot the astronaut with the drone. He was unable to see it on the live feed but was later able to view it clearly during post-processing. Hopefully, in a real-life situation, the astronaut doesn’t mind waiting a few extra hours for rescue… Meanwhile, the other three EVA members walked towards the ledge and found the transmitter lying exactly where we expected it to be, which was surprising given the high winds and rain from a Martian storm we experienced a few days ago. Unfortunately, we found the box halfway down the cliff facing Compass Rock. While we could have scrambled down the cliff to rescue it, the crew made an executive decision to put our safety first and to leave it behind. Sometimes we astronauts must follow the pirate’s code. After all, dead men tell no tales. With that, we gathered up the transmitter booty, took some epic pictures and drone videos, left the box to its fate in Davy Jones’ locker, and sailed back to the Hab.

As a fun addition to our mission, Space Kidz India sent us some great questions from K-12 students about life on Mars and we had a great time answering a few of them while on this EVA. We discussed topics including the Martian atmosphere and why spacesuits are necessary, as well as what a Mars analog mission is and why it is important (and fun!).

And with that, we concluded another successful EVA and ended Crew 272’s final in-sim expedition. Ever. New friends, incredible views, and an experience we’ll never forget. Thank you MDRS.

“I may not have gone where I intended to go, but I think I have ended up where I intended to be” – Douglas Adams.

Destination: Ledge overlooking Compass Rock (named Boiler Ledge by Crew 272)

Coordinates (use UTM WGS 84): N519600, E4251500

Participants: Megan Rush (HSO), Mason Kuhn (Crew Engineer), Arly Black (XO/Crew Scientist), Kenny Pritchard (Crew Journalist)

Road(s) and routes per MDRS Map: Drove north along Cow Dung Road until Gateway to Candor and parked the rovers. From Gateway to Candor, walked on foot to N519600, E4251500 and retrieved the transmitter. Returned on foot to the parked rovers and drove back to the Hab along Cow Dung Road.

Mode of travel: Walking and driving

Sol Summary Report – January 13th



Crew 272 Sol Summary Report 13-01-2023

Sol: 12

Summary Title: Rescuing the BTS: The Grand Finale

Author’s name: Kshitij Mall

Mission Status: All nominal

Sol Activity Summary: The grand finale of our Martian crew ensued with a brief morning warm up exercise session, yoga, and meditation led by me. The latter were needed to get the crew calm, composed, and focused as a lot of tasks were on our plate. But before anything else, there was something else that was on our morning breakfast plate: the biscuits purely prepared from Martian resources and hot coffee to wake up the crew and partially fill their veins with the necessary caffeine to keep the crew running. The breakfast for the sol was prepared by Crew Geologist Adriana and supported by Crew Journalist Kenneth. The humongous mission of this sol was to retrieve my electrical baby named the Boiler Transmission Station (BTS) and we named this mission Dynamite. The dynamic EVA crew for the Dynamite mission comprised Crew Engineer (and Pilot-in-Command) Mason, XO Arly, Kenneth, and HSO Megan. Kenneth was a last-minute entry as I had to delegate some other important tasks to myself at the Hab. A huge shoutout to the Purdue Mission Control back at Earth (especially Rachana and Alex) for providing us with the METAR data just in time to help Mason with decision making regarding the drone flight.

Upon leaving the front airlock of the Hab Sweet Hab, the EVA crew was in mission mode to find the BTS. Within 10 minutes they parked their Martian horses, Spirit and Oppie, next to the Gateway to Candor, and then started the transmitter hunt. Mason then launched his flying baby, Garud the drone, and started looking for the BTS. He used visual clues to locate the Boiler Ledge at first because that’s where we last saw BTS. The whole EVA crew slowly started marching up the hill next to the Galileo Road and toward the Boiler Ledge. Garud the drone had brief connectivity issues as encountered previously possibly due to terrain issues. The EVA crew spotted some aliens from a high point but ignored them as the aliens were departing Mars. Mason couldn’t see clearly on the live feed from Garud but thought of checking the video generated post-processing. The last-minute entry to the EVA crew, Kenneth spotted the transmitter eventually in association with Garud and proved to be the hero that I expected him to be while swapping spots with him for this mission. He performed well in the Geologist versus Non-Geologist experiments too as per Adriana and has seen BTS from a close proximity. The EVA crew then also found the BTS box halfway down the cliff, north of the Campus rock, which was expected due to strong winds 2 sols back. Since it was not safe to rescue the box, the crew settled for taking back just the transmitter. After the successful retrieval of the BTS (sans the box) and completion of Mission Dynamite, the crew took some drone media footage to celebrate the successful search and rescue operation. The EVA crew also took group pictures including some with the Purdue flag, which were apt for a place like the Boiler Ledge. Arly and Kenneth also recorded some videos for our Outreach activity with Space Kidz India, which we will be sending back in once we are out of sim and back to Earth. These questions were relayed to Space Kidz India team by several school students hailing from different parts of India. Our plan is to answer remaining outreach questions at the Hab and surrounding units. Even though I couldn’t go to rescue my electrical baby, I’m glad my fellow Martian crew mates rescued it and brought it back to the Hab. BTS was reunited with his sibling transmitter at the Hab, and it was an electrically emotional moment.

As per traditions, the crew completed all the cognitive performance tests and daily questionnaires requested by Arly and Andres. At the Greenhab, Madelyn watered the plants and will fertilize the soil in the afternoon after which she plans to clean this unit. The Science Dome will be deprived of the Martian rocks we collected so far as those rocks are being packed by Adriana in a big box to be taken back to Purdue for further analysis. We plan to clean the Hab and surrounding units and then to pack our Earthly and Martian belongings with a Divide and Conquer policy aimed at dividing the cleanup work among the 7 of us. Of course, we stay united while we do all these cleanup activities. The final mopping and vacuuming of the lower deck will be carried out tomorrow morning to keep the Hab cleanest possible when we depart for our journey back to the Earth. As I submit this final sol summary report for our rotation, we are already out of sim and ready to launch to our other sweet home (a Hab away from Hab). We will surely miss the Martian times but will be ready to savor back the Earthly flavors before we return back to Mars. We thank the Mission support team for an outstanding support to keep us safe, sound, and healthy during the rotation with replenishments and shipments via interplanetary cargo ships.

Look Ahead Plan: Sending the final research and daily (sol-ly) reports, packing our stuff, cleaning the Hab and surrounding units, and then relaxing a bit at MDRS

Anomalies in work: None

Weather: Overcast with Calm Winds

Crew Physical Status: Good

EVA: One

Reports to be filed: Sol Summary, Operations Report, Journalist Report, EVA Report, Greenhab Report, Mission Summary

Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site